Parth (parthkohli):

I had this question on my test. If \(\cot \theta + \tan \theta = x\) And \(\sec \theta - \cos \theta = y\) Then evaluate \(\left( x^2 y + xy^2\right)^{2/3}\) Also, could you guys just show me the solution, as I wouldn't be able to respond you with equations? A hint would also be fine, but it should be enough for me to know how to solve the whole problem. Thanks!

OpenStudy (anonymous):

cot x= 1/tanx and sec x = 1/cos x

Parth (parthkohli):

Aware of that. Let me try the problem again though, just to be sure I need you guys.

OpenStudy (anonymous):

ok if a post a solution then wud it be prob?

OpenStudy (anonymous):

write in terms of cos and sin ..will get answer pretty fast

OpenStudy (dls):

\[\LARGE x=\frac{\cos^2 \theta+\sin^2 \theta}{2\sin \frac{\theta}{2}}=\frac{1}{2} cosec \frac{\theta}{2}\] \[\LARGE y=\frac{1-\cos^2 \theta}{\cos \theta}=\frac{\sin^2 \theta}{\cos \theta}=\tan \theta \sec \theta\] \[(x^2y+xy^2)^\frac{2}{3}=(\frac{1}{4}cosec^2 \frac{\theta}{2} \tan \theta \sec \theta+\frac{1}{2}cosec \frac{\theta}{2}\tan^2 \theta \sec^2 \theta)^\frac{2}{3}\] hmmm..

Parth (parthkohli):

How should I work toward my answer? I tried that but didn't seem to work...

Parth (parthkohli):

Huh

OpenStudy (nincompoop):

lol

Parth (parthkohli):

Wut

OpenStudy (nincompoop):

I think it is plain substitution

OpenStudy (dls):

Tuw

Parth (parthkohli):

Wouldn't that make it lengthy?

OpenStudy (nincompoop):

maybe that's the purpose then to evaluate by simplification

Parth (parthkohli):

Yeah, but what to simplify?

OpenStudy (dls):

how much simplified form do u want? i did it :|

Parth (parthkohli):

The answer is a number

OpenStudy (dls):

:3

Parth (parthkohli):

Wait, did you do the complete answer or left the rest to me?

OpenStudy (dls):

left the rest to you

Parth (parthkohli):

A problem is that we haven't done half angle identities... only the three identities.

OpenStudy (dls):

xy(x+y) secx cosecx tanx sec x (secxcosecx+tanxsecx) sec^3x( secxcosecx+tanxsecx) \[\frac{1}{\cos x \sin x} (\frac{cosx}{\sin x \cos^3 x}+ \frac{\sin x \cos x}{\sin x \cos^3x})=\] \[\Large \frac{1}{\sin x}(\frac{1+\sin x}{\sin x \cos^3x})=\frac{1+\sin x}{\sin^2x \cos^3 x}\] \[\Large (\frac{1+\sin x}{\sin^2x \cos^3 x})^\frac{2}{3}\]=?

OpenStudy (nincompoop):

alternatively ((2 sin(theta)+cos(2 theta)-3)/(sin(2 theta)-2 cos(theta)))^(2/3)

OpenStudy (dls):

WF result^ :P

OpenStudy (dls):

I don't think we can get a numerical answer from this question anyway

OpenStudy (nincompoop):

if he said it is a number then it is a number.

OpenStudy (nincompoop):

ya it is a lot of work, so I just plugged it in wolfram

OpenStudy (nincompoop):

lazy in the AM syndrome

OpenStudy (usukidoll):

wolfram is awesome, but what happens if they have to do it manually meaning no computers?

OpenStudy (nincompoop):

screw manual then I do it in class all the time

OpenStudy (dls):

if you have to do manually then do easier one like mine :D

OpenStudy (usukidoll):

what about on an exam? I have some professors who don't allow calculators.

OpenStudy (goformit100):

2/2 = 1

OpenStudy (dls):

we have a winner here!

OpenStudy (nincompoop):

yeah. it is easy, just tedious

OpenStudy (usukidoll):

yeah I don't like tedious problems. I'm trying to memorize formulas here. better now than later since I'll be using them in future sections

OpenStudy (usukidoll):

I had some problems where I did all the computation for a one or a zero as an answer

OpenStudy (dls):

this wil happen many times in your life :) especially in limits

OpenStudy (usukidoll):

yay a medal!

OpenStudy (usukidoll):

in the other thread...didn't realize that

OpenStudy (nincompoop):